OpenStudy (anonymous):

find x and y intercepts of 3x-4y=-12

OpenStudy (anonymous):

put it in to y=mx+b form -4y=3-12 now solve for y y=-(3/4)x+3

OpenStudy (anonymous):

your y at x=0 it 3

OpenStudy (anonymous):

your slope is rise/run therefore -3/4 from (0,3) move down 3 and over 4 this will give you x at y=0 4

OpenStudy (anonymous):

y inter=3 x inter=4 with a negative slope

OpenStudy (anonymous):

thank you